# Some Density Results for HyperKähler Manifolds

### v1

@@ 2.1. Wall Divisors and the Kähler cones
Here, $\operatorname{Mon}^{2}(X)_{\rm Hdg}$ is the subgroup of monodromy operators that preserve the Hodge structure on $H^{2}(X,\mathbb{C})$, and $D^{\perp}=\{\alpha\in\overline{\mathcal{C}_{X}}\mid q_{X}(\alpha,D)=0\}.$ Thus a line bundle $L$ is nef on $X$ provided $q_{X}(L,D)\geq 0$ for all wall divisors $D\in\operatorname{Pic}(X).$
@@ 2.2. Proof of Theorem 4
Let $t\in\operatorname{Def}(X,L)\setminus\operatorname{Def}(X,L)_{\operatorname{nef}}$. Since $L_{t}\coloneqq\mathcal{L}|_{X_{t}}$ is not nef on $X_{t}$, there exists a wall divisor $D_{t}\in\operatorname{Pic}(X_{t})$ such that $D_{t}^{\perp}\cap\overline{\mathcal{K}_{X_{t}}}\neq\varnothing$, i.e., the wall $D_{t}^{\perp}$ is a wall of the Kähler cone, and $q_{X_{t}}(D_{t},L_{t})<0$. Then the class $D_{t}\in H_{2}(X,{\mathbb{Z}})$ generates an extremal ray of the Mori cone and some positive rational multiple $kD_{t}$ is represented by an effective rational curve $C\subset X_{t}.$
@@
Put $h:=\varphi_{t}(H)$, and denote by $\Omega_{h}\coloneqq\Omega\cap\mathbb{P}(h^{\perp}\otimes\mathbb{C})$, the associated period domain. We denote by $\operatorname{Mon}(\Lambda,h)\coloneqq\{g\in\operatorname{Mon}(X_{t})\mid g(h)=h\}$. Since the lattice $h^{\perp}\subset\Lambda$ has signature $(2,*)$, by [Huy16, Remark 6.1.10] the group $\operatorname{Mon}(\Lambda,h)$ acts properly discontinuously on $\Omega_{h}$ and hence for any $v\in h^{\perp}$, the union $\bigcup_{g\in\operatorname{Mon}(\Lambda,h)}g(v)^{\perp}$ is closed. By [AV20, Theorem 3.7], the monodromy group $\operatorname{Mon}(\Lambda,h)$ acts on the set of wall divisors ${\mathcal{W}}$ with finitely many orbits. Therefore the set
@@ Acknowledgments
We are grateful to BIRS in Banff, and to the organizers of the workshop “Women in Mathematical Physics II” in August 2023, where this project started. We would like to thank Ignacio Barros, Emma Brakkee, Daniel Huybrechts, and Claire Voisin for interesting conversations and comments on the results. The third named author is partially supported by a grant from the Simons Foundation/SFARI (522730, LK).

### v3

@@ 2.1. Wall Divisors and the Kähler cones
Here, $\operatorname{Mon}^{2}(X)_{\rm Hdg}$ is the subgroup of monodromy operators that preserve the Hodge structure on $H^{2}(X,\mathbb{C})$ (see [Mar11, Defn. 1.1]), and $D^{\perp}=\{\alpha\in\overline{\mathcal{C}_{X}}\mid q_{X}(\alpha,D)=0\}.$ Thus a line bundle $L$ is nef on $X$ provided $q_{X}(L,D)\geq 0$ for all wall divisors $D\in\operatorname{Pic}(X).$
@@ 2.2. Proof of Theorem 4
Let $t\in\operatorname{Def}(X,L)\setminus\operatorname{Def}(X,L)_{\operatorname{nef}}$. Since $L_{t}\coloneqq\mathcal{L}|_{X_{t}}$ is not nef on $X_{t}$, there exists a wall divisor $D_{t}\in\operatorname{Pic}(X_{t})$ such that $D_{t}^{\perp}\cap\overline{\mathcal{K}_{X_{t}}}\neq\varnothing$, i.e., the wall $D_{t}^{\perp}$ is a wall of the Kähler cone, and $q_{X_{t}}(D_{t},L_{t})<0$. Then, by [Mon15, Prop 1.5], the class $D_{t}\in H_{2}(X_{t},{\mathbb{Z}})$ generates an extremal ray of the Mori cone (the closed convex cone generated by the classes of effective curves), and some positive rational multiple $kD_{t}$ is represented by an effective curve $C\subset X_{t}.$ By the Cone Theorem, these effective curves are rational when $X_{t}$ is projective; in the non-projective case, it was observed in [AV20], that these extremal curces are the so-called MBM curves and are rational.
@@
Put $h:=\varphi_{t}(H)$, and denote by $\Omega_{h}\coloneqq\Omega\cap\mathbb{P}(h^{\perp}\otimes\mathbb{C})$, the associated period domain. We denote by $\operatorname{Mon}(\Lambda,h)\coloneqq\{g\in\operatorname{Mon}(X_{t})\mid g(h)=h\}$. Since the lattice $h^{\perp}\subset\Lambda$ has signature $(2,*)$, by [Huy16, Remark 6.1.10] the group $\operatorname{Mon}(\Lambda,h)$ acts properly discontinuously on $\Omega_{h}$ and hence for any $v\in h^{\perp}$, the union $\bigcup_{g\in\operatorname{Mon}(\Lambda,h)}g(v)^{\perp}$ is a proper closed subset. By [AV20, Theorem 3.7], the monodromy group $\operatorname{Mon}(\Lambda,h)$ acts on the set of wall divisors ${\mathcal{W}}$ with finitely many orbits. Therefore the set
@@ Acknowledgments
We are grateful to BIRS in Banff, and to the organizers of the workshop “Women in Mathematical Physics II” in August 2023, where this project started. We would like to thank Ignacio Barros, Emma Brakkee, Daniel Huybrechts, and Claire Voisin for interesting conversations and comments on the results. We thank Christian Lehn, Giovanni Mongardi, Salim Tayou, and the referee for their useful comments. The third named author is partially supported by a grant from the Simons Foundation/SFARI (522730, LK).